# Assessing HRV and HR Dynamics with Wearables During Socially Anxious Situations: Insights from a Controlled Study in a Low-Middle-Income Country

### v1

@@
Abstract.
This paper investigates physiological markers of Social Anxiety Disorder (SAD) by examining the relationship between Electrocardiogram (ECG) measurements and speech, a known anxiety-inducing activity. Specifically, we analyze changes in heart rate variability (HRV) and heart rate (HR) during four distinct phases: baseline, anticipation, speech activity, and reflection. Our study, involving 51 participants (31 with SAD and 20 without), found that HRV decreased and HR increased during the anticipation and speech activity phases compared to baseline. In contrast, during the reflection phase, HRV increased and HR decreased. Additionally, participants with SAD exhibited lower HRV, higher HR, and reported greater self-perceived anxiety compared to those without SAD. These findings have implications for developing wearable technology to monitor SAD. We also provide our dataset, which captures anxiety across multiple stages, to support further research in this area.
Social anxiety disorder, Physiological markers, Wearables
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Mathematics and statistics
1. Introduction
@@
Social Anxiety Disorder (SAD), one of the anxiety spectrum disorder, is characterized by an overwhelming sense of fear and apprehension in social settings. SAD, also known as Social Phobia, significantly impairs one’s ability to carry out daily tasks because of an intense fear of being negatively evaluated by others in various social situations, such as conversing with strangers or participating in classroom discussions. Consequently, individuals with SAD tend to avoid social situations due to the fear of embarrassment. As of 2020, around 19% of adults in the US population (approx. 40 million individuals) are affected by the anxiety disorder (lecamwasam2023investigating, ). Considering that anxiety ranks as one of the most prevalent mental health disorders worldwide, detecting and addressing it at a younger age is essential.
Conventionally, SAD is diagnosed through clinical diagnostic interviews or self-reported questionnaires. Clinical interviews involve a psychiatrist or counselor conducting an in-depth conversation with an individual afflicted by SAD. However, it is found that such individuals tend to avoid clinical visits due to their time-consuming nature, which often necessitates multiple appointments (kolaas2023feasibility, ). On the other hand, self-reported questionnaires require individuals to respond to a predefined set of questions, often using a Likert scale. However, self-report measures get hindered due to poor recall or bias, thus limiting their accuracy (akbar2019email, ). These limitations of the conventional diagnostic approaches result in the treatment gap and advocate for new, reliable, and scalable methods for the timely detection of SAD.
@@
The lack of reliable and consistent physiological markers for SAD highlights the need to examine variations in various physiological parameters — such as heart rate, saliva composition, and blood pressure — among individuals in different anxious social situations. Since these parameters are controlled by the Autonomic Nervous System (ANS), which is sensitive to anxiety-inducing stimuli (lecamwasam2023investigating, ), it is crucial to explore how they vary under different conditions and types of anxiety. Among the various physiological makers, Heart Rate (HR) and Heart Rate Variability (HRV) can be easily monitored nowadays as the wearables come with pre-installed Photoplethysmography (PPG) sensor. However, research on HR and HRV variability during anxiety-inducing activities, such as public speaking or cognitive tasks, shows mixed results. Some studies report “significantly reduced” HRV (alvares2013reduced, ; gaebler2013heart, ; garcia2017autonomic, ), while others report “insignificantly reduced or increased” HRV (held2021heart, ; pittig2013heart, ; tolin2021psychophysiological, ; cheng2022heart, ) during anxiety provoking situations in SAD individuals. Notably, Held et al. (held2021heart, ) report that these mixed outcomes are not necessarily due to the nature of anxiety-inducing activities but rather stem from the differing characteristics of participants. These variations in participant characteristics raise concerns about the global generalizability of existing research findings. Therefore, it becomes imperative to re-evaluate the relationship between cardiovascular measures (i.e., HR, HRV) and anxiety-provoking activities across different world regions.
@@
In this paper, we introduce a novel study that examines for the first time how HR and HRV fluctuate over time—specifically before, during, and after an anxiety-provoking activity—within a low-middle-income country (LMIC). Our approach is innovative because it separately analyzes HR and HRV variations across three distinct phases: Anticipation, activity, and reflection compared to the Baseline. This contrasts with the majority of existing research, which typically compares HR/HRV variations only at a single time point or immediately after an anxiety-inducing event (alvares2013reduced, ; miranda2014anxiety, ; tamura2013salivary, ; madison2021social, ).
@@
Our study significantly advances the field by contributing to the intersection of Human-Computer Interaction (HCI) research and mental health assessment by investigating potential physiological markers of SAD. Conducted in a controlled lab setting, the study involved ninety-nine non-clinical university students undergoing an anxiety-provoking task. Participants were assessed during the Baseline phase and across three additional phases—anticipation, activity, and reflection. We classified participants into SAD and non-SAD groups based on their Social Phobia Inventory (SPIN) scores, a standard measure for SAD severity. Figure 1 provides a high-level overview of our study. To test our following hypotheses, we employed hierarchical linear modeling (HLM), adding a novel methodological dimension to the analysis of anxiety-related physiological responses.
@@
  • Hypothesis 1: There will be a significant increase in HR and a decrease in HRV during the Anticipation and Speech activity phases, respectively, compared to the Baseline for all participants. Additionally, HR and HRV will exhibit significant differences during the Reflection phase compared to the Baseline phase for all participants. This hypothesis does not consider grouping participants based on different attributes (e.g., gender, age) and relies solely on HR and HRV analysis.
@@
  • Hypothesis 2: There will be significant differences in HR and HRV between the SAD and non-SAD participants during the anxious activity. This hypothesis considers grouping participants based on their anxiety severity (i.e., SAD, Non-SAD).
@@
  • Hypothesis 3: There will be a significant increase in perceived anxiety levels during the Anticipation and Speech activity phases compared to the Baseline for all participants. Furthermore, perceived anxiety levels during the Reflection phase will significantly differ from those during the Baseline. This hypothesis differs from the previous two hypotheses as it investigates “perceived anxiety” instead of HR or HRV.
@@
From our hypotheses testing, we have determined significant patterns in the participants’ physiological responses. Specifically, we observed reduced HRV during the anticipation and activity phases compared to the baseline, while HRV increased during the reflection phase for all participants. Moreover, we observed that SAD individuals exhibited lower HRV compared to non-SAD. Our analysis revealed that HRV parameters such as RMSSD [^1] The square root of the mean of squared successive differences between adjacent RR intervals, SD1222The standard deviation perpendicular to the identity line, reflecting short-term RR interval fluctuations, i.e., beat-to-beat variability, and SDNN [^3] The standard deviation of RR intervals can discriminate between participants with SAD and those without it. Our examination of perceived anxiety levels indicated that all participants experienced heightened anxiety during both the anticipation and activity phases compared to the baseline. However, anxiety levels decreased during the reflection phase, highlighting a dynamic interplay between physiological responses and subjective emotional experiences across different study phases. Again, we found that SAD individuals exhibited higher perceived anxiety levels compared to non-SAD.
@@
Figure 1. Each study session involved a Research Associate (RA) and three participants, labeled P1, P2, and P3, represented by pink, blue, and yellow colors. Initially, ECG sensors were attached to the participants, and they provided self-reported SPIN responses. The participants then proceeded through four phases: Baseline, Anticipation, Speech Activity, and Reflection. After the Baseline phase and each subsequent phase, participants reported their perceived anxiety levels (PAL) using a single-item questionnaire. A total of 40 study sessions were conducted. For data analysis, HR and HRV features—covering time-domain, frequency-domain, and non-linear metrics—were extracted from cleaned ECG signals and analyzed using a multilevel model to determine their potential as physiological markers for SAD.
Following are the contributions of this paper.
  1. (1)
    We contribute to the fields of HCI and affective computing by designing a novel three-phase study to investigate changes in the HR, HRV, and perceived anxiety during anticipation, activity, and reflection phases, separately. The study design and the generated insights are unique from the existing literature. One of our unique findings is that participants perceive higher anxiety during anticipation compared to the actual anxious activity phase, showing anticipation is more anxiety-provoking that the actual activity. This finding can be used as a guideline while designing future HCI and affective computing studies to understand the relationship between anxious activities, perceived anxiety, and physiological markers.
  2. (2)
    We compile a unique dataset from a controlled study on SAD (our-hrv-dataset, ). This dataset includes HR, HRV metrics ( time domain, frequency domain, and nonlinear dynamics), and self-reported anxiety levels. Its uniqueness lies in its study design, which captures both objective measures (HR and HRV) and self-reported anxiety at three different stages of an anxiety-provoking activity. Additionally, our dataset includes participants’ SPIN scores and demographic information (gender, age, location).
    To our knowledge, there are no publicly available HRV datasets on SAD that offer such comprehensive information. This dataset can serve as a valuable benchmark for researchers studying anxiety disorders, as it includes data from both SAD and non-SAD participants. It is also useful for exploring the relationships between gender, age, location, self-reported scores, HRV, and HR. To promote collaboration and advance research in mental health, we are making this dataset publicly available alongside this paper.
  3. (3)
    Our study presents findings from a LMIC. Cultural differences significantly impact how anxiety is experienced and expressed (vaillant2012positive, ; eshun2009introduction, ). Existing research (fernando2019developing, ; salamanca2020ethical, ; de2015missed, ; asher2017little, ) indicates that solutions for anxiety may need to be adapted to fit regional contexts, as the symptoms and underlying causes can vary widely. By contributing data from an LMIC, our study adds valuable perspectives to the existing body of research, primarily from developed countries. To our knowledge, no other study from an LMIC has investigated SAD in controlled settings, whether clinical or non-clinical.
  4. (4)
    We rigorously test the above-stated hypotheses using hierarchical linear models and performed a correlation analysis of several important features. Hierarchical models provide a powerful framework for analyzing data with complex structures, offering more nuanced insights and improving the accuracy and interpretability of statistical inferences. Our analysis identifies important features for the HCI design community.
  5. (5)
    Additionally, our work contributes to affective computing for mental health assessment by using physiological changes to recognize and interpret human emotions. This research aligns with the CHI community’s mission to design, develop, and evaluate innovative technologies that enhance human life. By focusing on the physiological changes that occur before, during, and after an anxious activity, our study offers critical insights into the embodied experience of anxiety. These physiological markers not only provide a more accurate and reliable means of assessing anxiety levels but also serve as potential targets for future interventions, paving the way for personalized, real-time mental health support systems.
Our work is structured as follows: Section 2 presents related work on study design in controlled settings for anxiety physiological marker detection and the use of wearable sensors to detect mental disorders. Section 3 details our methodology, including participant recruitment and the study procedures we followed. Section 4 discusses the analysis we performed to test our hypothesis. Section 5 outlines our findings and their implications. Sections 5.6 and 5.7 address our work’s limitations and suggest future research directions. Finally, Section 6 provides concluding remarks.
2. Literature Review
@@
SAD has attracted attention from various academic disciplines, including computer science, physiology, psychology, and psychiatry, resulting in a diverse body of literature. In the context of our research, we have categorized related works into four distinct sections: (i) Mental Health and HCI: This section sheds light on research by HCI researchers on mental health, (ii) Study Designs: This section classifies existing research examining cardio-vascular measures of SAD based on their study designs, i.e., distinguishing between single-phase and multiple-phase studies. (iii) Anxiety-Inducing Activities: This section sheds light on the various anxiety-inducing activities employed across different studies to investigate SAD. (iv) Wearables and smartphones for mental health disorder detection: This section discusses the role of wearables and smartphones in controlling mental health disorders. The details of each of the mentioned sections are below.
2.1. Mental Health and HCI
In recent years, a growing body of research has emerged at the intersection of affective computing and digital mental health, driven by collaborations between mental health professionals and HCI researchers. HCI has been at the forefront of exploring technological solutions to understand and address mental health issues. This body of work primarily focuses on developing practical solutions for predicting (cai2023cathill, ; alslaity2022insights, ; kim2022prediction, ; kruzan2022wanted, ; lecamwasam2023investigating, ; nepal2024moodcapture, ) or controlling (hamid2022you, ; baghaei2020time, ; zhang2021designing, ; lattie2020designing, ; lecamwasam2023investigating, ; kruzan2023perceived, ; hwang2024societal, ) mental health conditions in innovative ways. For instance, Nepal et al. (nepal2024moodcapture, ) utilized smartphone cameras to capture daily images of participants, aiming to predict depression severity. Similarly, Lecamwasam et al. (lecamwasam2023investigating, ) investigated the impact of music interventions on physiological signals during stress-inducing activities. These studies have significantly contributed to the development of both group-level and personalized interventions in the field of affective computing for mental health. Our proposed research continues this trajectory by aiming to discover physiological markers of SAD through a controlled study involving participants doing an anxious activity.
2.2. Study designs
@@
Research shows that individuals suffering from SAD may exhibit cardiovascular responses that are different from those without the disorder (lecamwasam2023investigating, ; held2021heart, ). HR and HRV are the most commonly measured cardiovascular physiological responses. HR measures the number of heartbeats per minute, while HRV quantifies the variations in time intervals between heartbeats, offering insights into how the ANS adapts to various stimuli (havard_hrv, ; held2021heart, ). Based on the study design, existing SAD studies can be categorized as single-phase or multiple-phase, as described below.
@@
Similarly, Miranda et al. (miranda2014anxiety, ) and Tamura et al. (tamura2013salivary, ) analyzed HR and HRV during a presentation and electrical stimulation, respectively, and did not observe any significant increase in HR among SAD participants. Furthermore, Tamura et al. (tamura2013salivary, ) reported no significant differences in the frequency domain HRV parameters (LF, HF, LF/HF) between SAD and non-SAD participants.
ii. Multiple-phase studies: These studies investigate dynamic changes in HR and HRV before, during, and after engaging in anxiety-inducing activity. They provide a comprehensive view of how these parameters fluctuate in response to anxiety. For example, Tolin et al. (tolin2021psychophysiological, ) and Rubio et al. (garcia2017autonomic, ) examined changes in HRV across the baseline, Trier Social Stress Test (TSST) activity, and the recovery phase. However, the findings from these studies were different, i.e., Rubio et al. (garcia2017autonomic, ) observed that SAD participants show significantly reduced RMSSD, while Tolin et al. (tolin2021psychophysiological, ) did not identify any significant changes in RMSSD. Nevertheless, both studies reported a common finding: SAD participants exhibited higher HR.
The HR and HRV findings vary across studies where some studies report “significantly reduced” HRV (alvares2013reduced, ; gaebler2013heart, ; garcia2017autonomic, ), while others report “insignificantly reduced or increased” HRV (held2021heart, ; pittig2013heart, ; tolin2021psychophysiological, ; cheng2022heart, ) in SAD participants. Similarly, most (miranda2014anxiety, ; tamura2013salivary, ; gaebler2013heart, ) studies did not find increased HR, whereas Alavares et al.(alvares2013reduced, ) found Higher HR in SAD participants. The mixed findings necessitate further investigation to comprehensively understand the role of HRV and HR as potential physiological markers for SAD.
@@
2.3. Anxiety-inducing activities
In controlled settings, often anxiety-inducing activities are used to understand the effect of an anxious activity on different physiological (e.g., ECG) and psychological responses of participants. These activities encompass a wide range of tasks, including speech (harrewijn2018heart, ; garcia2017autonomic, ; grossman2001gender, ), hyperventilation (pittig2013heart, ), cognitive challenges (held2021heart, ; gaebler2013heart, ; tolin2021psychophysiological, ; garcia2017autonomic, ; madison2021social, ), reading (garcia2017autonomic, ), presentations (miranda2014anxiety, ), and behavioral assessment tests such as opposite-sex interaction (asher2020dating, ; asher2020out, ). Table 1 lists various anxiety-inducing activities employed in previous studies. It is worth noting that the duration of these anxiety-inducing activities varies across studies, reflecting the diverse approaches used to explore the physiological and psychological responses of participants with SAD in different contexts.
2.4. Wearables and smartphones for mental disorder detection
Over the past decade, research has increasingly explored the use of wearables (xue2022understanding, ; lecamwasam2023investigating, ) and smartphones (salekin2018weakly, ; rashid2020predicting, ; wang2023detecting, ) for detecting physiological and behavioral markers of mental disorders. For instance, Lecamwasam et al. (lecamwasam2023investigating, ) employed Empatica E4 sensors to monitor participant’s HR and Electrodermal activity (EDA) during mathematical problem-solving tasks, accompanied by personalized music to examine its impact on stress and anxiety induced by the tasks. Likewise, Xue et al. (xue2022understanding, ) used chest-mounted ECG heart rate sensors and explored how group members collectively reflected on organizational stress using a shared and anonymous visualization of HRV. Notably, most studies involving wearables have primarily focused on stress (wang2022first, ; adler2021identifying, ; mishra2020evaluating, ; yu2023semi, ; xue2022understanding, ; lecamwasam2023investigating, ).
Although the symptoms of mental disorders are consistent globally, their manifestations and underlying causes can differ due to societal and cultural variations (meegahapola2023generalization, ). Thus, it is important to conduct studies in diverse geographical locations to uncover the unique causes of anxiety in different regions. In this context, our proposed wearable-based controlled study is a pioneering effort from a lower-middle-income country to identify potential physiological markers of SAD. To our knowledge, most wearable-based studies on anxiety have been conducted in developed countries. We believe our findings will complement and extend the insights gained from research in developed nations.
Table 1. Summary of SAD studies studying variations in HR and HRV features during different anxious activities.

### v2

# Exploring Heart Rate Variability and Heart Rate Dynamics Using Wearables Before, During, and After Speech Activity: Insights from a Controlled Study in a Low-Middle-Income Country

@@
Abstract
Conventional methods for diagnosing Social Anxiety Disorder (SAD), such as clinical interviews and self-reported questionnaires, often face accessibility barriers and subjective biases, underscoring the need for objective physiological markers. This study investigates heart rate (HR) and heart rate variability (HRV) as potential indicators of SAD by analyzing cardiovascular responses to anxiety-inducing speech tasks across four distinct phases: baseline, anticipation, speech activity, and reflection.
In a controlled laboratory setting, we analyzed data from 51 participants and found that HRV decreased and HR increased during the anticipation and speech activity phases compared to baseline, while the reflection phase showed a reversal, with HRV increasing and HR decreasing. Participants with SAD exhibited lower HRV, higher HR, and greater self-reported anxiety than non-SAD participants across all phases. These findings enhance our understanding of the physiological signatures of social anxiety and have implications for developing wearable-based monitoring systems for SAD detection and intervention. To support further research, we also release a dataset capturing multi-phase anxiety responses, advancing physiological-based mental health assessment. [^1] This is the author’s accepted manuscript of a paper accepted for publication in ACM COMPASS. The final published version is available via the ACM Digital Library at <https://doi.org/10.1145/3715335.3735477>..
1 Introduction
@@
Social Anxiety Disorder (SAD), one of the anxiety spectrum disorders, is characterized by an overwhelming sense of fear and apprehension in social settings. SAD, also known as Social Phobia, significantly impairs one’s ability to carry out daily tasks because of an intense fear of being negatively evaluated by others in various social situations, such as conversing with strangers or participating in classroom discussions. Consequently, individuals with SAD tend to avoid social situations due to the fear of embarrassment. As of 2020, around 19% of adults in the US population (approx. 40 million individuals) are affected by the anxiety disorder [lecamwasam2023investigating]. Considering that anxiety ranks as one of the most prevalent mental health disorders worldwide, detecting and addressing it at a younger age is essential.
@@
The lack of reliable and consistent physiological markers for SAD highlights the need to examine variations in various physiological parameters — such as heart rate, saliva composition, and blood pressure — among individuals in different anxious social situations. Since these parameters are controlled by the Autonomic Nervous System (ANS), which is sensitive to anxiety-inducing stimuli [lecamwasam2023investigating], it is crucial to explore how they vary under different conditions and types of anxiety. Among the various physiological makers, Heart Rate (HR) and Heart Rate Variability (HRV) can be easily monitored nowadays as the wearables come with pre-installed Photoplethysmography (PPG) sensor. However, research on HR and HRV variability during anxiety-inducing activities, such as public speaking or cognitive tasks, shows mixed results. Some studies report “significantly reduced” HRV [alvares2013reduced, gaebler2013heart, garcia2017autonomic], while others report “insignificantly reduced or increased” HRV [held2021heart, pittig2013heart, tolin2021psychophysiological, cheng2022heart] during anxiety provoking situations in SAD individuals. Notably, Held et al. [held2021heart] report that these mixed outcomes are not necessarily due to the nature of anxiety-inducing activities but rather stem from the differing characteristics of participants. These variations in participant characteristics raise concerns about the global generalizability of existing research findings. Therefore, it becomes imperative to re-evaluate the relationship between cardiovascular measures (i.e., HR, HRV) and anxiety-provoking activities across different world regions.
@@
In this paper, we introduce a novel study that examines for the first time how HR and HRV fluctuate over time—specifically before, during, and after an anxiety-provoking activity—within a low-middle-income country (LMIC). Our approach is innovative because it separately analyzes HR and HRV variations across three distinct phases: Anticipation, Activity, and Reflection compared to the Baseline. This contrasts with the majority of existing research, which typically compares HR/HRV variations only at a single time point or immediately after an anxiety-inducing event [alvares2013reduced, miranda2014anxiety, tamura2013salivary, madison2021social].
Our study makes an important contribution to affective computing, focusing on the Global South—a region that faces unique and often overlooked challenges in mental health. To our knowledge, this is the first three-phase study that explores the use of physiological markers for identifying mental health disorders in India. In LMICs such as India, mental health issues are shaped by complex, intersecting factors such as social stigma, limited access to care, and cultural perceptions of illness [pendse2019mental]. These factors deeply shape both the experience and treatment of mental health conditions, often exacerbating the suffering of affected individuals [patel2018lancet].
@@
We conducted a controlled lab study involving ninety-nine non-clinical university students who underwent an anxiety-inducing task. Participants were assessed during the Baseline phase and across three additional phases—Anticipation, Activity, and Reflection. We classified participants into SAD and non-SAD groups based on their Social Phobia Inventory (SPIN) scores, a standard measure for SAD severity. Figure 1 provides a high-level overview of our study. To test our following hypotheses, we employed hierarchical linear modeling (HLM), adding a novel methodological dimension to the analysis of anxiety-related physiological responses.
@@
  • Hypothesis 1 (H1): There will be a significant increase in HR and a decrease in HRV during the Anticipation and Activity phases, respectively, compared to the Baseline for all participants. Additionally, HR and HRV will exhibit significant differences during the Reflection phase compared to the Baseline phase for all participants. This hypothesis does not consider grouping participants based on different attributes (e.g., gender, age) and relies solely on HR and HRV analysis.
@@
  • Hypothesis 2 (H2): There will be significant differences in HR and HRV between the SAD and non-SAD participants during the Activity phase. This hypothesis considers grouping participants based on their anxiety severity (i.e., SAD, Non-SAD).
@@
  • Hypothesis 3 (H3): There will be a significant increase in perceived anxiety levels (self-reported) during the Anticipation and Activity phases compared to the Baseline for all participants. Furthermore, perceived anxiety levels during the Reflection phase will significantly differ from those during the Baseline. This hypothesis differs from the previous two hypotheses as it investigates “perceived anxiety” instead of HR or HRV.
@@
From our hypotheses testing, we have determined significant patterns in the participants’ physiological responses. Specifically, we observed reduced HRV during the anticipation and activity phases compared to the baseline, while HRV increased during the reflection phase for all participants. Moreover, we observed that SAD participants exhibited lower HRV compared to non-SAD. Our analysis revealed that HRV parameters such as RMSSD [^2] The square root of the mean of squared successive differences between adjacent RR intervals, SD1333The standard deviation perpendicular to the identity line, reflecting short-term RR interval fluctuations, i.e., beat-to-beat variability, S [^4] The area of an ellipse described by SD1 and SD2, proportional to SD1SD2, and SDNN [^5] The standard deviation of RR intervals can discriminate between participants with SAD and those without it. Our examination of perceived anxiety levels indicated that all participants experienced heightened anxiety during both the anticipation and activity phases compared to the baseline. However, anxiety levels decreased during the reflection phase, highlighting a dynamic interplay between physiological responses and subjective emotional experiences across different study phases. Again, we found that SAD participants exhibited higher perceived anxiety levels compared to non-SAD. Following are the contributions of this paper.
@@
Figure 1: Each study session involved a Research Associate (RA) and three participants, labeled P1, P2, and P3, represented by pink, blue, and yellow colors. Initially, ECG sensors were attached to the participants, and they provided self-reported SPIN responses. The participants then proceeded through four phases: Baseline, Anticipation, Speech Activity, and Reflection. After the Baseline phase and each subsequent phase, participants reported their perceived anxiety levels (PAL) using a single-item questionnaire. For data analysis, HR and HRV features—covering time-domain, frequency-domain, and non-linear metrics—were extracted from cleaned ECG signals and analyzed using a multilevel model to determine their potential as physiological markers for SAD.
  • We design and implement novel three-phase study to investigate changes in the HR, HRV, and perceived anxiety during anticipation, activity, and reflection phases, separately. Unlike prior work, we analyze each phase separately to capture temporal dynamics of anxiety.
  • We compile and publicly release a unique, high-resolution dataset [^6] https://osf.io/phze7/ from a controlled laboratory study on SAD conducted in India. To the best of our knowledge, this is the first publicly available HRV dataset focused on SAD in a Global South context.
  • Our study contributes a culturally grounded understanding of anxiety in the Global South [pendse2019mental]. Cultural factors significantly shape how anxiety is experienced and expressed [vaillant2012positive, eshun2009introduction]. Existing literature [fernando2019developing, salamanca2020ethical, de2015missed, asher2017little] underscores the need for region-specific mental health research. To our knowledge, this is the first controlled study of SAD conducted in Indian setting.
2 Literature Review
@@
SAD has attracted attention from various academic disciplines, including computer science, physiology, psychology, and psychiatry, resulting in a diverse body of literature. In the context of our research, we have categorized related works into three distinct sections: (i) Study designs: This section classifies existing research examining cardio-vascular measures of SAD based on their study designs, i.e., distinguishing between single-phase and multiple-phase studies. (ii) Anxiety-inducing activities: This section sheds light on the various anxiety-inducing activities employed across different studies to investigate SAD. (iii) Wearables and smartphones for mental health disorder detection: This section discusses the role of wearables and smartphones in controlling mental health disorders. The details of each of the mentioned sections are below.
2.1 Study designs
@@
Research shows that individuals suffering from SAD may exhibit cardiovascular responses that are different from those without the disorder [lecamwasam2023investigating, held2021heart]. HR and HRV are the most commonly measured cardiovascular physiological responses. HR measures the number of heartbeats per minute, while HRV quantifies the variations in time intervals between heartbeats, offering insights into how the ANS adapts to various stimuli [havard_hrv, held2021heart]. Based on the study design, existing SAD studies can be categorized as single-phase or multiple-phase, as described below.
@@
Similarly, Miranda et al. [miranda2014anxiety] and Tamura et al. [tamura2013salivary] analyzed HR and HRV during a presentation and electrical stimulation, respectively, and did not observe any significant increase in HR among SAD participants. Furthermore, Tamura et al. [tamura2013salivary] reported no significant differences in the frequency domain HRV parameters (LF, HF, LF/HF) between SAD and non-SAD participants.
ii. Multiple-phase studies: These studies investigate dynamic changes in HR and HRV before, during, and after engaging in anxiety-inducing activity. They provide a comprehensive view of how these parameters fluctuate in response to anxiety. For example, Tolin et al. [tolin2021psychophysiological] and Rubio et al. [garcia2017autonomic] examined changes in HRV across the baseline, Trier Social Stress Test (TSST) activity, and the recovery phase. However, the findings from these studies were different, i.e., Rubio et al. [garcia2017autonomic] observed that SAD participants show significantly reduced RMSSD, while Tolin et al. [tolin2021psychophysiological] did not identify any significant changes in RMSSD. Nevertheless, both studies reported a common finding: SAD participants exhibited higher HR.
The HR and HRV findings vary across studies where some studies report “significantly reduced” HRV [alvares2013reduced, gaebler2013heart, garcia2017autonomic], while others report “insignificantly reduced or increased” HRV [held2021heart, pittig2013heart, tolin2021psychophysiological, cheng2022heart] in SAD participants. Similarly, most [miranda2014anxiety, tamura2013salivary, gaebler2013heart] studies did not find increased HR, whereas Alavares et al.[alvares2013reduced] found Higher HR in SAD participants. The mixed findings necessitate further investigation to comprehensively understand the role of HRV and HR as potential physiological markers for SAD.
@@
2.2 Anxiety-inducing activities
In controlled settings, often anxiety-inducing activities are used to understand the effect of an anxious activity on different physiological (e.g., ECG) and psychological responses of participants. These activities encompass a wide range of tasks, including speech [harrewijn2018heart, garcia2017autonomic, grossman2001gender, sahu2024unveiling], hyperventilation [pittig2013heart], cognitive challenges [held2021heart, gaebler2013heart, tolin2021psychophysiological, garcia2017autonomic, madison2021social], reading [garcia2017autonomic], presentations [miranda2014anxiety], and behavioral assessment tests such as opposite-sex interaction [asher2020dating, asher2020out]. Table 1 lists various anxiety-inducing activities employed in previous studies. It is worth noting that the duration of these anxiety-inducing activities varies across studies, reflecting the diverse approaches used to explore the physiological and psychological responses of participants with SAD in different contexts.
2.3 Wearables and smartphones for mental disorder detection
Over the past decade, research has increasingly explored the use of wearables [xue2022understanding, lecamwasam2023investigating] and smartphones [salekin2018weakly, rashid2020predicting, wang2023detecting] for detecting physiological and behavioral markers of mental disorders. For instance, Lecamwasam et al. [lecamwasam2023investigating] employed Empatica E4 sensors to monitor participant’s HR and Electrodermal activity (EDA) during mathematical problem-solving tasks, accompanied by personalized music to examine its impact on stress and anxiety induced by the tasks. Likewise, Xue et al. [xue2022understanding] used chest-mounted ECG heart rate sensors and explored how group members collectively reflected on organizational stress using a shared and anonymous visualization of HRV. Notably, most studies involving wearables have primarily focused on stress [wang2022first, adler2021identifying, mishra2020evaluating, yu2023semi, xue2022understanding, lecamwasam2023investigating].
Although the symptoms of mental disorders are consistent globally, their manifestations and underlying causes can differ due to societal and cultural variations [meegahapola2023generalization]. Thus, it is important to conduct studies in diverse geographical locations to uncover the unique causes of anxiety in different regions. In this context, our proposed wearable-based controlled study is a pioneering effort from a lower-middle-income country to identify potential physiological markers of SAD. To our knowledge, most wearable-based studies on anxiety have been conducted in developed countries. We believe our findings will complement and extend the insights gained from research in developed nations.
Table 1: Summary of SAD studies studying variations in HR and HRV features during different anxious activities.